# Exploring the mysterious high-ionization source powering [NeĀ V] in high- analog SBS0335-052Ā E with JWST/MIRI

### v1

@@ Abstract
Nearby blue compact dwarf galaxies (BCDs) share similar properties with objects from the Epoch of Reionization revealed by JWST, in terms of low stellar mass, low metallicity and high specific star-formation rate. Thus, they represent ideal local laboratories for detailed multi-wavelength studies to understand their properties and the mechanisms shaping them. We report the first JWST MIRI/MRS observations of the BCD SBSĀ 0335-052Ā E, analyzing MIR emission lines tracing different levels of ionization (e.g., [NeĀ II], [SĀ IV], [NeĀ III], [OĀ IV], [NeĀ V]) of the ionized gas. SBSĀ 0335-052Ā E MIR emission is characterized by a bright point source, located in one of the youngest and most embedded stellar clusters (Ā Myr, ), and underlying extended high-ionization emission (i.e., [OĀ IV], [NeĀ V]) from the surroundings of the older and less dusty stellar clusters (Ā Myr, ). From the comparison with state-of-the-art models, we can exclude shocks, X-ray binaries, and old stellar populations as the main sources of the high ionization. Interestingly, a 4-16% contribution of a Ā Mā intermediate massive black hole (IMBH) is needed to justify the strong [NeĀ V]/[NeĀ II]Ā and would be consistent with optical/UV line ratios from previous studies. However, even IMBH models cannot explain the strongest [OĀ IV]/[NeĀ III]. Also, star-forming models (regardless of including X-ray binaries) struggle to reproduce even the lower ionization line ratios (e.g., [SĀ IV]/[NeĀ II]) typically observed in BCDs. Overall, while current models suggest the need to account for an accreting IMBH in this high- analog, limitations still exist in predicting high-ionization emission lines (I.P.Ā Ā eV) when modeling these low-metallicity environments, thus other sources of ionization cannot be fully ruled out.
@@ 1 Introduction
In this context, nearby galaxies can provide a precious opportunity for us to understand the characteristics of the interstellar medium (ISM) of these primeval objects and how to distinguish between star-formation and AGN activity within them, given the high signal-to-noise (S/N) of their spectra and the possibility to perform spatially resolved and multi-wavelength studies. In particular, nearby blue compact dwarf galaxies (BCDs) are ideal analogs to these high- systems, as they share properties in terms of stellar mass, intense star formation activity and metal poor content, as also suggested by recent JWST studies (e.g., Schaerer etĀ al. 2022; Brinchmann 2023). Many studies have explored these objects to understand how to constrain their ISM properties using their optical and UV emission (e.g., James etĀ al. 2009; Izotov etĀ al. 2012; Berg etĀ al. 2019; 2021; Senchyna etĀ al. 2021; 2023; Mingozzi etĀ al. 2022; 2024; Kumari etĀ al. 2024). In this context, the mid-infrared (MIR) wavelength range covers a large range of ionization potentials for the same species (e.g., 20-126Ā eV, [NeĀ II]Ā 21.56m, [NeĀ III]Ā 15.56m, [NeĀ V]Ā 14.32,24.3m, [NeĀ VI]Ā ) at lower excitation energies. While these lines are not accessible in the high- universe, they can provide an important constraint on the gas ionization mechanisms and, by also probing other phases (i.e., warm molecular gas with H2 rotational lines; dust with polycyclic aromatic hydrocarbons, PAHs), they can give a broad overview of the ISM conditions.
@@
In this work, we focus on the galaxy SBSĀ 0335-052Ā E (; 1āĀ Ā pc), discovered by Izotov etĀ al. (1990). This object is a nearby BCD, extremely metal-poor (Ā Zā; 12+log(O/H)Ā ; Papaderos etĀ al. 2006) and characterized by a compact vigorous starburst (specific star formation rate, log(/Gyr-1)Ā ; RĆ©my-Ruyer etĀ al. 2015) and massive young clusters (Papaderos etĀ al. 1998; Reines etĀ al. 2008; Adamo etĀ al. 2010 and references therein). This galaxy is also well-known for a complex kinematics and a large scale HĀ outflow (Herenz etĀ al. 2023) and high-ionization optical and UV emission lines typically observed in the EoR (e.g., HeĀ IIĀ 4686, CĀ IVĀ , HeĀ IIĀ 1640; Herenz etĀ al. 2017; Kehrig etĀ al. 2018; Wofford etĀ al. 2021). In particular, its extended nebular HeĀ IIĀ 4686 emission requires energy beyond 54.4Ā eV (i.e., Ā Ryd, Ā Ć ), indicative of a very hard ionizing spectral energy distribution (SED), that only rotating metal-free or extremely metal-poor massive stars seem capable of reproducing (Kehrig etĀ al. 2018; see also Wofford etĀ al. 2021). Recent studies have also suggested that SBSĀ 0335-052Ā E may host an intermediate massive black hole (IMBH) of Ā Mā to explain its [NeĀ V]Ā 3426 emission (Hatano etĀ al. 2024, see also Thuan & Izotov 2005), requiring energies beyond 97Ā eV (i.e., Ā Ryd, Ā Ć ), as well as possible near IR variability (Hatano etĀ al. 2023). IMBHs (see Greene etĀ al. 2020 for a review) could represent the missing piece between stellar and supermassive BHs and would help explain the rapid accretion processes in place at high-. Thus, it is important to understand if galaxies such as SBSĀ 0335-052Ā E can host an accreting IMBH, and, if so, which are the best diagnostics to trace its activity and impact on the ISM.
@@
To this end, we exploit JWST/MIRI Medium Resolution Spectroscopy (MRS) integral field spectroscopy (IFS) data to investigate the possible accreting IMBH or other non-stellar source ionizing the high- analog SBSĀ 0335-052Ā Eās ISM, using spatially resolved high ionization MIR lines. Indeed, SBS 0335-052Ā E intense SF activity makes this galaxy particularly bright in the MIR, which represents % of its total luminosity (Plante & Sauvage 2002). Spitzer IRS Ā m spectrum already revealed strong MIR emission lines (i.e., [SĀ IV]Ā 10.51, [NeĀ III]Ā 15.55, [OĀ IV]Ā 25.89), warm molecular gas features (i.e., H2Ā Ā S(3)) and the absence of PAHs (Houck etĀ al. 2004; Hao etĀ al. 2009). Nevertheless, MIRI/MRS offer the considerable advantage of spatially resolving the MIR emission (down to ā, Ā pc) and thus revealing the location of very high ionization extended emission (i.e., [NeĀ V]Ā 14.32m) for the first time in BCDs.
@@
In Sec.Ā 2, we present the observations and data reduction. In Sec.Ā 3, we show the methods we adopted for our data analysis, including spectral fitting and literature photoionization and shock models that we used to interpret our results, reported in Sec.Ā 4. In Sec.Ā 5, we discuss our findings as well as possible limitations in current models, while in Sec.Ā 6, we conclude with a summary of our work. Throughout this work we assume a flat CDM cosmology (HĀ km/s/Mpc, ) and 12+log(O/H) (Asplund etĀ al. 2009).
@@ 2 Observations and Data Reduction
SBS 0335-052Ā E data were acquired as part of the Cycle 2 JWST program PID 4278 (PI: Mingozzi) on February 11, 2024. In particular, the galaxy was observed with JWST/MIRI MRS using the three MRS grating settings (SHORT, MEDIUM, and LONG; Ā ), and the FAST readout mode. Just one pointing was used since MIRI/MRS covers the majority of SBSĀ 0335-052Ā Eās active starburst activity with ChannelĀ 1 (3.2ā3.7ā) and fully samples it in ChannelĀ 4 (6.6ā7.7ā). Given the JWST pointing accuracy of ā, target acquisition was not required. A 4-point dither pattern was also applied to achieve optimal sampling throughout the MRS field of view (FOV) and to identify and remove detector artefacts. Each pointing was observed with 50 groups per integration for a total on-source exposure time of 1665Ā s. Dedicated background observations were performed with half the integrations (same group number but half the dithering) to accurately measure and correct for the thermal background. Background and science observations were connected via a fixed offset and placed in an uninterruptible sequence, and a PA orientation range was required to obtain simultaneous imaging observations using the F770W, F1130W, and F1800W filters, which will be presented in a forthcoming paper.
@@
We retrieved MIRI/MRS raw products from Mikulski Archive for Space Telescopes (MAST) and we processed the data with v1.16.0 of the JWST calibration pipeline, following the jupyter notebook provided by the MIRI team. As extra steps to the standard version of the MIRI pipeline, we masked additional regions affected by cosmic showers at the end of StageĀ 1 (ShowerMasking Notebook) and we applied residual fringe corrections in StageĀ 2. We subtracted the background in StageĀ 3, using the master background created by the pipeline. In StageĀ 3 we created the final datacubes per band, to optimize the fringing correction, obtaining 12 datacubes centered to the same coordinates and keeping the native pixel scale size (i.e., 0.13ā, 0.17ā, 0.20ā and 0.35ā, for Channel 1, 2, 3 and 4, respectively). Nevertheless, we note that few spaxels located at the peak of the MIR emission are still affected by significant fringing, especially in ChannelsĀ 3 and 4. The FOV progressively increases from 3.2ā3.7ā in Channel 1, to 6.6ā7.7ā in Channel 4, as well as the MIRI/MRS PSF and spatial resolution, from in Channel 1 to in Channel 4.
@@
In Fig.Ā 1 we show the position of MIRI/MRS FOV superimposed on the HST ACS F150LP UV emission (left panel) and the 12-17Ā m slice from MIRI/MRS Channel 3 (right panel). The UV emission (map and grey contours) clearly shows the six young, massive and compact āsuper star clustersā (SSC1, SSC2, SSC3, SSC4, SSC5, SSC6) revealed in this galaxy (Thuan etĀ al. 1997). SSCs1-6 clusters have ages and masses in the range Ā Myr and Ā Mā (Reines etĀ al. 2008; Adamo etĀ al. 2010; see also Tab.Ā 2), and luminosity of Ā Lā, all within āĀ Ā pc (Thuan etĀ al. 1997). One of them, SSC1, is also known for its enhanced obscuration (Hunt etĀ al. 2001). The HeĀ IIĀ 4686 high-ionization emission revealed with MUSE (white dotted contours; seeing ā) follows the UV emission and peaks around SSCs4,5. The HST H black dashed contours highlight the three additional actively young star-forming regions, S3-Pa, S7 and S8, discovered by Thompson etĀ al. (2009) by analyzing SBSĀ 0335-052Ā E Pa emission. The MIRI/MRS image (right panel) reveals the clear MIRI/MRS PSF pattern due to a very bright point source located at the position of SSC1 as well as extended emission covering the other stellar clusters, as we will discuss in Sec.Ā 3.
@@ 3.1 Emission-line fitting
In each spectrum we fitted several MIR emission lines, performing the line fitting locally around each line in a km/s window centered at the observed wavelength of the line of interest. In particular, we used a combination of a Gaussian profile and a one-degree polynomial to account for emission-line and surrounding continuum emission in each spaxel of the point-source subtracted band datacube slices, as well as in the 1-D spectrum of the bright point-source and the spectrum extracted at the ULX position (shown in Fig.Ā 1). The 1-D spectrum of the bright point source is obtained by the best-fit model of the PSF photometry analysis described in App.Ā A, while the ULX spectrum is extracted from a region with variable diameter depending on the channel, matching the FWHM of the PSF (i.e., Ā 0.4ā, 0.5ā, 0.6ā, 0.9ā (Law etĀ al. 2023). Table 1 compiles the list of emission lines analyzed, together with their rest-frame wavelengths and ionization potentials (I.P.).
@@
Fig.Ā 2, 3 and 4 show the point source spectrum (SSC1), the spectrum extracted from the spaxels covering the ULX position, and a spectrum from the north-west region covering the clusters SSCs3,4,5,6 and S7 and the peak of UV emission (see Fig.Ā 1), respectively, as well as the best-fit (in red) for the MIR emission lines. In the following, we consider line detections as those with fluxes with signal-to-noise (S/N) higher than 3 and upper limits for measurements with S/N2. In particular, we note that emission lines typically observed in AGN (e.g., Dasyra etĀ al. 2024), such as [NeĀ VI]Ā Ā 7.65 (tentatively fitted only in the bright point source spectrum with ), [MgĀ V]Ā Ā 5.61 (I.P.Ā Ā eV) and [MgĀ VII]Ā Ā 5.5 (I.P.Ā Ā eV) are not detected.
@@ 3.2 Dust attenuation
The emission-line fluxes were converted from the original MJy/sr into erg/s/cm2/arcsec2 and corrected for the total foreground Galactic reddening along the line of sight using the values determined for this object by Berg etĀ al. (2022) and the Gordon etĀ al. (2023) dust extinction relationship.
@@
When exploring the spectrum of the point source SSC1, we noticed a wide absorption in the 7.5-12.5m range, corresponding to the silicate feature at 9.7Ā m (highlighted in grey in Fig.Ā 2, 3 and 4). This feature (revealed also in SBSĀ 0335-052Ā E Spitzer spectrum; Houck etĀ al. 2004) pinpoints the presence of dust within this region, as already reported by Hunt etĀ al. (2014). Thus, an intrinsic dust attenuation correction has to be applied to the line fluxes as well. Before using the continuum to determine the level of dust attenuation, there are few extra corrections we needed to implement to the datacubes. First, we limited the impact of fringing on the continuum spectra by applying an additional 1D fringing correction using the JWST pipeline tools. Also, the data exhibit discontinuities, or ājumpsā, in the flux normalization among the different bands, which we corrected by fixing the normalization to Band A in Channel 1. With the continuum corrected and optimized for the analysis, we used the Continuum and Feature Extraction software (CAFE, Marshall etĀ al. 2007; Diaz-Santos et al., in prep.) to model the continuum of both the point-source spectrum and the point-source subtracted (extended) continuum in the range Ā m from Channel 1, 2 and 3 band datacubes (see App.Ā A). CAFE uses a combination of blackbody emission at different temperatures together with a custom attenuation law (compatible with Gordon etĀ al. 2023) to describe the overall shape of the continuum at near-to-mid infrared wavelengths. For the case of SBS 0335-052Ā E only two dust components were considered: a cool component, whose temperature falls in the range 40-100 K, and a warm component, with temperature in the range 100-500 K. While both spectra, the point-like source and the extended emission, needed a cool component of 80 K, only the point-like source required an additional warm component with 150 K to model the continuum shape. A more detailed description of the dust attenuation correction will be presented in del Valle-Espinosa et al., in prep. As a result, the peak of dust attenuation corresponds to the point source SSC1 with , while for the extended region we considered an average value of . We used the derived values and the Gordon etĀ al. (2023) dust extinction relationship to correct for intrinsic dust attenuation.
@@
Although the extinction curve at the MIR wavelengths is almost flat, lines located in the deepest part of the silicate feature (highlighted in grey in Fig.Ā 2, 3 and 4) can suffer from large attenuation. In the case of SBSĀ 0335-052Ā E, the dust-corrected fluxes of almost every emission line are increased by a factor of () in the point-like source (extended) spectrum, except for the [ArĀ III]Ā and [SĀ IV]Ā lines ā which fall within the silicate feature ā whose correction factor is (). This also implies that only emission-line ratios including [ArĀ III]Ā and [SĀ IV]Ā lines can be affected by dust attenuation.

### v2

@@ Abstract
Nearby blue compact dwarf galaxies (BCDs) are considered analogs to objects from the Epoch of Reionization revealed by JWST, having similar low stellar mass, low metallicity, and high specific star-formation rate. Thus, they represent ideal local laboratories detailed for multi-wavelength studies of their properties and mechanisms that shape them. We report the first JWST MIRI/MRS observations of the BCD SBSĀ 0335-052Ā E, analyzing MIR emission lines tracing different levels of ionization (e.g., [NeĀ II], [SĀ IV], [NeĀ III], [OĀ IV], [NeĀ V]) of the ionized gas. SBSĀ 0335-052Ā E MIR emission is characterized by a bright point source, located in one of the youngest and most embedded stellar clusters (Ā Myr, ), and underlying extended high-ionization emission (i.e., [OĀ IV], [NeĀ V]) from the surroundings of the older and less dusty stellar clusters (Ā Myr, ). From the comparison with state-of-the-art models, we can exclude shocks, X-ray binaries, and old stellar populations as the main sources of ionization. Interestingly, a 4-8% contribution of a Ā Mā intermediate massive black hole (IMBH) is needed to justify the strong [NeĀ V]/[NeĀ II]Ā and would be consistent with optical/UV line ratios from previous studies. However, even IMBH models cannot explain the strongest [OĀ IV]/[NeĀ III]. Also, star-forming models (regardless of including X-ray binaries) struggle to reproduce even the lower ionization line ratios (e.g., [SĀ IV]/[NeĀ II]) typically observed in BCDs. Overall, while current models suggest the need to account for an accreting IMBH in this high- analog, limitations still exist in predicting high-ionization emission lines (I.P.Ā Ā eV) when modeling these low-metallicity environments, thus other sources of ionization cannot be fully ruled out.
@@ 1 Introduction
In this context, nearby galaxies can provide a precious opportunity for us to understand the characteristics of the interstellar medium (ISM) of these primeval objects and how to distinguish between star-formation and AGN activity within them, given the high signal-to-noise (S/N) of their spectra and the possibility to perform spatially resolved and multi-wavelength studies. In particular, nearby blue compact dwarf galaxies (BCDs) are ideal analogs to these high- systems, as they share similar properties in terms of stellar mass, intense star formation activity, and poor metal content, as also suggested by recent JWST studies (e.g., Schaerer etĀ al. 2022; Brinchmann 2023). Many studies have explored these objects to understand how to constrain their ISM properties using their optical and UV emission (e.g., James etĀ al. 2009; Izotov etĀ al. 2012; Berg etĀ al. 2019; 2021; Senchyna etĀ al. 2021; 2023; Mingozzi etĀ al. 2022; 2024; Kumari etĀ al. 2024). In this context, the mid-infrared (MIR) wavelength range covers a large range of ionization potentials for the same species (e.g., 20-126Ā eV, [NeĀ II]Ā 21.56m, [NeĀ III]Ā 15.56m, [NeĀ V]Ā 14.32,24.3m, [NeĀ VI]Ā 7.65m) at lower excitation energies. While these lines are not accessible in the high- universe, they can provide an important constraint on the gas ionization mechanisms and, by also probing other phases (i.e., warm molecular gas with H2 rotational lines; dust with polycyclic aromatic hydrocarbons, PAHs), they can give a broad overview of the ISM conditions.
@@
In this work, we focus on the galaxy SBSĀ 0335-052Ā E (; 1āĀ Ā pc), discovered by Izotov etĀ al. (1990). This object is a nearby BCD, extremely metal-poor (Ā Zā; 12+log(O/H)Ā ; Papaderos etĀ al. 2006) and characterized by a compact vigorous starburst (specific star formation rate, log(/Gyr-1)Ā ; RĆ©my-Ruyer etĀ al. 2015) and massive young clusters (Papaderos etĀ al. 1998; Reines etĀ al. 2008; Adamo etĀ al. 2010 and references therein). This galaxy is also well-known for complex kinematics and a large scale HĀ outflow (Herenz etĀ al. 2023) and high-ionization optical and UV emission lines typically observed in the EoR (e.g., HeĀ IIĀ 4686, CĀ IVĀ , HeĀ IIĀ 1640; Herenz etĀ al. 2017; Kehrig etĀ al. 2018; Wofford etĀ al. 2021). In particular, its extended nebular HeĀ IIĀ 4686 emission requires energy beyond 54.4Ā eV (i.e., Ā Ryd, Ā Ć ), indicative of a very hard ionizing spectral energy distribution (SED), that only rotating metal-free or extremely metal-poor massive stars seem capable of reproducing (Kehrig etĀ al. 2018; see also Wofford etĀ al. 2021). Recent studies have also suggested that SBSĀ 0335-052Ā E may host an intermediate massive black hole (IMBH) of Ā Mā to explain its [NeĀ V]Ā 3426 emission (Hatano etĀ al. 2024, see also Thuan & Izotov 2005), requiring energies beyond 97Ā eV (i.e., Ā Ryd, Ā Ć ), as well as possible near IR variability (Hatano etĀ al. 2023). IMBHs (see Greene etĀ al. 2020 for a review) could represent the missing piece between stellar and supermassive BHs and would help explain the rapid accretion processes in place at high-. Thus, it is important to understand if galaxies such as SBSĀ 0335-052Ā E can host an accreting IMBH, and, if so, which are the best diagnostics to trace its activity and impact on the ISM.
@@
To this end, we exploit JWST/MIRI Medium Resolution Spectroscopy (MRS) integral field spectroscopy (IFS) data to investigate the possible accreting IMBH or other non-stellar source ionizing the high- analog SBSĀ 0335-052Ā Eās ISM, using spatially resolved high ionization MIR lines. Indeed, SBS 0335-052Ā E intense SF activity makes this galaxy particularly bright in the MIR, which represents % of its total luminosity (Plante & Sauvage 2002). Spitzer IRS Ā m spectrum already revealed strong MIR emission lines (i.e., [SĀ IV]Ā 10.51, [NeĀ III]Ā 15.55, [OĀ IV]Ā 25.89), warm molecular gas features (i.e., H2Ā Ā S(3)) and the absence of PAHs (Houck etĀ al. 2004; Hao etĀ al. 2009). Nevertheless, MIRI/MRS offers the considerable advantage of spatially resolving the MIR emission (down to ā, Ā pc) and thus revealing the location of very high ionization extended emission (i.e., [NeĀ V]Ā 14.32m) for the first time in BCDs.
@@
In Sec.Ā 2, we present the observations and data reduction. In Sec.Ā 3, we show the methods we adopted for our data analysis, including spectral fitting, and literature photoionization and shock models that we used to interpret our results, reported in Sec.Ā 4. In Sec.Ā 5, we discuss our findings as well as possible limitations in current models, while in Sec.Ā 6, we conclude with a summary of our work. Throughout this work we assume a flat CDM cosmology (HĀ km/s/Mpc, ) and 12+log(O/H) (Asplund etĀ al. 2009).
@@ 2 Observations and Data Reduction
SBS 0335-052Ā E data were acquired as part of the Cycle 2 JWST program PID 4278 (PI: Mingozzi) on February 11, 2024. In particular, the galaxy was observed with JWST/MIRI MRS using the three MRS grating settings (SHORT, MEDIUM, and LONG; Ā ), and the FAST readout mode. Just one pointing was used since MIRI/MRS covers the majority of SBSĀ 0335-052Ā Eās active starburst activity with ChannelĀ 1 (3.2ā3.7ā) and fully samples it in ChannelĀ 4 (6.6ā7.7ā). Given the JWST pointing accuracy of ā, target acquisition was not required. A 4-point dither pattern was also applied to achieve optimal sampling throughout the MRS field of view (FOV) and to identify and remove detector artefacts. Each pointing was observed with 50 groups per integration for a total on-source exposure time of 1665Ā s. Dedicated background observations were performed with half the integrations (same group number but half the dithering) to accurately measure and correct for the thermal background. Background and science observations were connected via a fixed offset and placed in an uninterruptible sequence, and a PA orientation range was required to obtain simultaneous imaging observations using the F770W, F1130W, and F1800W filters, which will be presented in a future publication.
@@
We retrieved MIRI/MRS raw products from Mikulski Archive for Space Telescopes (MAST), and we processed the data with v1.16.0 of the JWST calibration pipeline (Bushouse etĀ al. 2022), following the jupyter notebook provided by the MIRI team. As extra steps to the standard version of the MIRI pipeline, following the recommendations from the JWST pipeline reduction team, we masked additional regions affected by cosmic showers at the end of StageĀ 1 (ShowerMasking Notebook) and we applied residual fringe corrections in StageĀ 2. We subtracted the background in StageĀ 3, using the master background created by the pipeline. In StageĀ 3, we created the final datacubes per band, to optimize the fringing correction, obtaining 12 datacubes centered to the same coordinates and keeping the native pixel scale size (i.e., 0.13ā, 0.17ā, 0.20ā and 0.35ā, for Channel 1, 2, 3 and 4, respectively). Nevertheless, we note that the spaxels surrounding the bright point source that we further discuss below, within a circular region with a radius equal to the MIRI/MRS PSF FWHM, are still significantly affected by fringing, especially in ChannelsĀ 3 and 4. The FOV progressively increases from 3.2ā3.7ā in Channel 1, to 6.6ā7.7ā in Channel 4, as well as the MIRI/MRS PSF and spatial resolution, from in Channel 1 to in Channel 4.
@@
In Fig.Ā 1, we show the position of MIRI/MRS FOV superimposed on the HST ACS F150LP UV emission (left panel) and the 12-17Ā m slice from MIRI/MRS Channel 3 (right panel). The UV emission (map on the left panel, gray contours on the right panel) clearly shows the six young, massive and compact āsuper star clustersā (SSC1, SSC2, SSC3, SSC4, SSC5, SSC6) revealed in this galaxy (Thuan etĀ al. 1997). SSCs1-6 clusters have ages and masses in the range Ā Myr and Ā Mā (Reines etĀ al. 2008; Adamo etĀ al. 2010; see also Tab.Ā 2), and luminosity of Ā Lā, all within āĀ Ā pc (Thuan etĀ al. 1997). One of them, SSC1, is also known for its enhanced obscuration (Hunt etĀ al. 2001). The HeĀ IIĀ 4686 high-ionization emission revealed with MUSE (white dotted contour on the right panel; seeing ā) follows the UV emission and peaks around SSCs4,5. The HST H black dashed contours highlight the three additional actively young star-forming regions, S3-Pa, S7 and S8, discovered by Thompson etĀ al. (2009) by analyzing SBSĀ 0335-052Ā E Pa emission. The MIRI/MRS image (right panel) reveals the clear MIRI/MRS PSF pattern due to a very bright point source located at the position of SSC1 as well as extended emission covering the other stellar clusters, as we will discuss in Sec.Ā 3.
@@ 3.1 Emission-line fitting
In each spectrum, we fitted several MIR emission lines, performing the line fitting locally around each line in a km/s window centered at the observed wavelength of the line of interest. In particular, we used a combination of a Gaussian profile and a one-degree polynomial to account for emission-line and surrounding continuum emission in each spaxel of the point-source subtracted band datacube slices, as well as in the 1-D spectrum of the bright point-source and the spectrum extracted at the ULX position (shown in Fig.Ā 1). The 1-D spectrum of the bright point source is obtained by the best-fit model of the PSF photometry analysis described in App.Ā A, while the ULX spectrum is extracted from a region with variable diameter depending on the channel, matching the FWHM of the PSF (i.e., Ā 0.4ā, 0.5ā, 0.6ā, 0.9ā; Law etĀ al. 2023). Tab.Ā 1 compiles the list of emission lines analyzed, together with their rest-frame wavelengths and ionization potentials (I.P.).
@@
Fig.Ā 2, 3 and 4 show the point source spectrum (SSC1), the spectrum extracted from the spaxels covering the ULX position, and a spectrum from the north-west region covering the older clusters SSCs3,4,5,6 and S7, the peak of UV emission and the ULX (see Fig.Ā 1), respectively, as well as the best-fit (in red) for the MIR emission lines. In the following, we consider line detections as those with fluxes with signal-to-noise (S/N) higher than 3 and upper limits for measurements with S/N2. In particular, we note that emission lines typically observed in AGN (e.g., Dasyra etĀ al. 2024), such as [NeĀ VI]Ā Ā 7.65 (tentatively fitted only in the bright point source spectrum but with ), [MgĀ V]Ā Ā 5.61 (I.P.Ā Ā eV) and [MgĀ VII]Ā Ā 5.5 (I.P.Ā Ā eV) are not detected.
@@ 3.2 Dust attenuation
The emission-line fluxes were converted from the original MJy/sr into erg/s/cm2/arcsec2 and corrected for the total foreground Galactic reddening along the line of sight using the values determined for this object by Berg etĀ al. (2022) and the Gordon etĀ al. (2023) dust extinction relationship (see also Gordon etĀ al. 2009; 2021; Decleir etĀ al. 2022; Fitzpatrick etĀ al. 2019).
@@
When exploring the spectrum of the point source SSC1, we noticed a wide absorption in the 7.5-12.5Ā m range, corresponding to the silicate feature at 9.7Ā m (highlighted in gray in Fig.Ā 2, 3 and 4). This feature (revealed also in SBSĀ 0335-052Ā E Spitzer spectrum; Houck etĀ al. 2004) pinpoints the presence of dust within this region, as already reported by Hunt etĀ al. (2014). Thus, an intrinsic dust attenuation correction has to be applied to the line fluxes as well. Before using the continuum to determine the level of dust attenuation, there are a few extra corrections we needed to implement to the datacubes. First, we limited the impact of fringing on the continuum spectra by applying an additional 1D fringing correction using the JWST pipeline tools. Also, the data exhibit discontinuities, or ājumpsā, in the flux normalization among the different bands, which we corrected by fixing the normalization to Band A in Channel 1. With the continuum corrected and optimized for the analysis, we used the Continuum and Feature Extraction software (CAFE, Marshall etĀ al. 2007; Diaz-Santos etĀ al. 2025) to model the continuum of both the point-source spectrum and the point-source subtracted (extended) continuum in the range Ā m from Channel 1, 2 and 3 band datacubes (see App.Ā A). CAFE uses a combination of blackbody emission at different temperatures together with a custom attenuation law (compatible with Gordon etĀ al. 2023 law; see also Gordon etĀ al. 2009; 2021; Decleir etĀ al. 2022; Fitzpatrick etĀ al. 2019) to describe the overall shape of the continuum at near-to-mid infrared wavelengths. For the case of SBS 0335-052Ā E only two dust components were considered: a cool component, whose temperature falls in the range 40-100 K, and a warm component, with temperature in the range 100-500 K. While both spectra, the point-like source and the extended emission, needed a cool component of 80 K, only the point-like source required an additional warm component with 150 K to model the continuum shape. A more detailed description of the dust attenuation correction will be presented in del Valle-Espinosa et al., in prep. As a result, the peak of dust attenuation corresponds to the point source SSC1 with , while for the extended region we considered an average value of . We used the derived values and the Gordon etĀ al. (2023) dust extinction relationship (see also Gordon etĀ al. 2009; 2021; Decleir etĀ al. 2022; Fitzpatrick etĀ al. 2019) to correct for intrinsic dust attenuation.
@@
Although the extinction curve at the MIR wavelengths is almost flat, lines located in the deepest part of the silicate feature (highlighted in gray in Fig.Ā 2, 3 and 4) can suffer from large attenuation. In the case of SBSĀ 0335-052Ā E, the dust-corrected fluxes of almost every emission line are increased by a factor of () in the point-like source (extended) spectrum, except for the [ArĀ III]Ā and [SĀ IV]Ā lines ā which fall within the silicate feature ā whose correction factor is (). This also implies that only emission-line ratios including [ArĀ III]Ā and [SĀ IV]Ā lines can be affected by dust attenuation.